OpenStudy (howard-wolowitz):

11. Assume that a sample is used to estimate a population proportion p. Find the margin of error E that corresponds to the given statistics and confidence level. Round the margin of error to four decimal places. 99% confidence; n = 6500, x = 1950 0.0128 0.0146 0.0083 0.011

OpenStudy (howard-wolowitz):

@kirbykirby

OpenStudy (howard-wolowitz):

someone told me this one is D is that true

OpenStudy (anonymous):

this is a test right if it is please try figuring out on your own if not I can help

OpenStudy (howard-wolowitz):

did i say it was?

OpenStudy (kirbykirby):

Hm well The margin of error is the part we saw before that's after the \(\pm\), so \[z\cdot \sqrt{\frac{p(1-p)}{n}}\]. And well for a 99% confidence level, z=2.58

OpenStudy (anonymous):

I didn't say u did did I I was asking u if it is a test !!!!!!!!!!!!!!!!!!!!!!

OpenStudy (howard-wolowitz):

its not

OpenStudy (howard-wolowitz):

what p

OpenStudy (anonymous):

I said I didn't say it was a test did I

OpenStudy (howard-wolowitz):

can u type the numbers where they are supposed to be then I"ll solve it

OpenStudy (kirbykirby):

um well \(p=\dfrac{x}{n}\) and \(z=2.58\) which is the z-score

OpenStudy (howard-wolowitz):

ok i see one sec

OpenStudy (howard-wolowitz):

wat number is p?

OpenStudy (kirbykirby):

\(p=\dfrac{x}{n}\), they give you x=1950, and n = 6500 :)

OpenStudy (howard-wolowitz):

then wat goes in for P?

OpenStudy (howard-wolowitz):

ok I got 0.0146

OpenStudy (kirbykirby):

that's what I got too

OpenStudy (howard-wolowitz):

12. Provide an appropriate response. Samples of size n = 900 are randomly selected from the population of numbers (0 through 9) produced by a random-number generator, and the proportion of numbers ≥ 5 is found for each sample. What is the distribution of the sample proportions? skewed to the right skewed to the left not enough information provided normal (approximately)

OpenStudy (howard-wolowitz):

\[\ge5\]

OpenStudy (howard-wolowitz):

usually these are normal

OpenStudy (kirbykirby):

yes it should be approx normal

OpenStudy (howard-wolowitz):

13. Solve the problem. The following confidence interval is obtained for a population proportion, p: (0.707, 0.745). Use these confidence interval limits to find the margin of error, E. 0.017 0.020 0.038 0.019

OpenStudy (kirbykirby):

So I explain the margin of error previously and you know the formula for the confidence interval. So.. you can represent the confidence interval as \(C.I = p\pm E\)

OpenStudy (kirbykirby):

The lower limit of the confidence interval would be \(p - E\) the upper limit would be \(p+E\) So: \(0.707=p-E\) \(0.745=p+E\)

OpenStudy (kirbykirby):

so basically it's a system of equations with 2 unknowns. You just need to solve for E :)

OpenStudy (howard-wolowitz):

0.019

OpenStudy (howard-wolowitz):

@kirbykirby

OpenStudy (howard-wolowitz):

@mathmate

OpenStudy (howard-wolowitz):

is this right

OpenStudy (kirbykirby):

ya I got that =]

OpenStudy (howard-wolowitz):

last one ok